OpenStudy (_malice_):

When x = 3 and y = 5, by how much does the value of 3x2 – 2y exceed the value of 2x2 – 3y ? Individual Question 4 14 16 20 50

OpenStudy (_malice_):

@tigerlover @T-Dawg02

OpenStudy (_malice_):

can someone give me a medal, I need 1 and then I can get a thing

OpenStudy (t-dawg02):

At x=3 and y=5: 3x^2-2y = 27-10 = 17 2x^2-3y = 18-15 = 3 17-3 = 14

OpenStudy (_malice_):

Could you explain in further detail

OpenStudy (welshfella):

you plug x = 3 and y = 5 into 3x^2 - 2y and work it out 3(3)^2 - 2(5) = 3*9 - 10 = 27-10 = 17

OpenStudy (t-dawg02):

When the value of x is equal to 3 and the y is equal to 5, 3x^2 – 2y => 3*3^2 - 2*5 => 3*9 - 10 = 17 And 2x^2– 3y => 2*3^2 - 3*5 => 2* 9 - 15 = 3 This gives 3x^2 – 2y exceeding 2x^2– 3y by 17 - 3 = 14

OpenStudy (_malice_):

Oooh, thank you very much :)

OpenStudy (_malice_):

I can't give yo both a medal sadly

OpenStudy (_malice_):

you*

OpenStudy (t-dawg02):

Give it to who ever you want

OpenStudy (_malice_):

I have to decide between a cat and a dog

OpenStudy (_malice_):

kek

OpenStudy (_malice_):

Uh oh, math war incoming

OpenStudy (_malice_):

I gtg but thanks for the help.

OpenStudy (t-dawg02):

Np